### v1

# Two-person positive shortest path games have Nash equlibria in pure stationary strategies

Abstract
We prove that every finite two-person positive shortest path game has a Nash equilibrium (NE) in pure stationary strategies, which can be computed in polynomial time. The existence result holds also for graphs with finite out-degrees. Moreover, we prove that a terminal NE exists provided at least one of two players can guarantee reaching a terminal. If no one can do it, in other words, if each of two players can cut all terminals from the initial position $s$, then, obviously, a cyclic NE exists, although its cost is infinite for both players, since we restrict ourselves to positive games. We conjecture that a terminal NE exists too, provided there exists a directed path from $s$ to a terminal. However, this is open.
Keywords: two-person shortest path games, Nash equilibrium in pure stationary strategies.
1 Introduction
1.1 Bi-shortest path theorem and conjecture
Let $\Gamma=(V,E)$ be a finite directed graph (digraph) that may have loops, multiple edges, and directed cycles. A fixed vertex $s\in V$ is called initial. Furthermore, $V$ is partitioned into $n+1$ subsets $V=V_{1},\cup\dots\cup V_{n}\cup V_{T}$ such that $E$ contains no edge $(v,w)$ leaving $v$ if and only if $v\in V_{T}$. In the latter case $v$ is called terminal or a terminal vertex.
Let us assign to each edge $e\in E$ $n$ real numbers $\{r_{i}(e)\mid i\in[n]=\{1,\dots,n\}\}$ and call $r_{i}(e)$ the $i$th local cost (or length) of $e$. Given $i\in[n]$ and a directed path $p$ from $s$ to a terminal $v\in V_{T}$, we assign the real number $R_{i}(p)=\sum_{e\in p}r_{i}(e)$ and call it the $i$th effective cost (or length) of $p$.
Denote by $\sigma_{i}:V_{i}\rightarrow V$ a mapping that assigns to each vertex $v\in V_{i}$ a vertex $w=\sigma_{i}(w)\in V$ such that $(v,w)\in E$ is a directed edge of $\Gamma$, and let $S_{i}$ be the set of such mappings.
Assume that all local costs are positive [^1] Equivalently, all directed cycles have positive total costs for each player [5]. ($r_{i}(e)>0$ for all $i\in[n]$ and $e\in E$) and set $n=2$ (that is, $[n]=\{1,2\}$). Fix a mapping $\sigma_{1}\in S_{1}$ and for each $v\in V_{1}$, choose the edge $(v,\sigma_{1}(v))\in E$ and delete from $\Gamma$ all other edges $(v,w)$ going from $v$. Denote the obtained reduced digraph $\Gamma(\sigma_{1})$ and denote by $P_{2}(\sigma_{1})$ the set of all shortest paths (SP) from $s$ to $V_{T}$ with respect to the local lengths $r_{2}$, or in other words, minimizing the effective cost $R_{2}$.
It may happen that $\sigma_{1}$ cuts $V_{T}$ from $s$, that is, digraph $\Gamma(\sigma_{1})$ has no directed path from $s$ to $V_{T}$ ($(s,V_{T})$-path). In this case we set $P_{2}(\sigma_{1})=\overline{\infty}$, where $\overline{\infty}$ is a special symbol, denoting a path of effective cost +$\infty$+ ∞. Set $P_{2}=\cup_{\sigma_{1}\in S_{1}}P_{2}(\sigma_{1})$. Then, by swapping 1 and 2 define $P_{1}$.
The main result of this paper, called the Bi-SP Theorem, is as follows.
Theorem 1
$P_{1}\cap P_{2}\neq\emptyset$.
In other words, sets $P_{1}$ and $P_{2}$ always have a directed path or symbol $\overline{\infty}$ in common. By definition, ¯$\overline{\infty}\in P_{i}$ for $i=1,2$ if and only if there exists a mapping $\sigma_{3-i}$ cutting $V_{T}$ from $s$, that is, the reduced digraph in $\Gamma(\sigma_{3-i})$ has no $(s,V_{T})$ path. In particular, ¯$\overline{\infty}\in P_{1}\cap P_{2}$ if and only if such cutting mappings $\sigma_{1}$ and $\sigma_{2}$ both exist. We conjecture that even in this case $P_{1}$ and $P_{2}$ have an $(s,V_{T})$-path in common provided it exists in $\Gamma$.
Conjecture 2
([3, 8, 9]) If $\Gamma$ has an $(s,V_{T})$ path then sets $P_{1}$ and $P_{2}$ have such a path in common, that is, $(P_{1}\cap P_{2})\setminus\{\overline{\infty}\}\neq\emptyset$.
Note that the case when $\Gamma$ has no $(s,V_{T})$ path is trivial. Indeed, in this case all considered sets contain only $\infty$ and no other elements.
As we know, ¯$\overline{\infty}\not\in P_{i}$ if and only if no $\sigma_{3-i}\in S_{i}$ cuts $V_{T}$ from $s$. If this happens for $i=1$ or $i=2$ then the Bi-SP Theorem and Conjecture coincide. Thus, the latter remains open only if there exist both $\sigma_{1}\in S_{1}$ and $\sigma_{2}\in S_{2}$ cutting $V_{T}$ from $s$. Otherwise, the statement of Conjecture 2 follows from Theorem 1.
[…]
1.2 Simplifying assumptions
Without loss of generality (wlog) we can assume that:
  • (j)
    $|V_{T}|=1$. Indeed, merge all terminal vertices in $V_{T}$ and replace them by a unique terminal $t$. After this, a directed path from $s$ to $V_{T}$ will be called an $(s,t)$-path.
  • (jj)
    An $(s,t)$-path exists. Indeed, otherwise all considered sets consist of a unique element $\overline{\infty}$.
  • (jjj)
    Each vertex $v\in V$ and edge $e\in E$ is reachable from $s$. Indeed, otherwise it can be just deleted.
The next two simplifying assumptions also can be made without loss of generality and, although are not needed for further proofs, but perhaps help the reader to understand the problem.
  • (jv)
    All directed $(s,t)$-paths have pairwise distinct effective costs. Indeed, we can get rid of all ties by an arbitrary small perturbation of local costs. After this, a shortest $(s,t)$-path is unique in $\Gamma(\sigma_{i})$, that is, $|P_{i}(\sigma_{3-i})|=1$ for all $i=1,2$. Actually, the same can be done for any $n$, not only for $n=2$.
  • (v)
    Digraph $\Gamma$ is bipartite, moreover, $u,w\not\in V_{i}$ for every edge $(u,w)\in E$ and $i\in[n]=\{1,2\}$. (The latter statement, holds, in fact, for any $n>1$, not only for $n=2$.) Indeed, let $(u,w)$ be such an edge. Then, subdivide it into two, that is, replace $(u,w)$ by two edges $(u,v)$ and $(v,w)$, where $v\in V_{j}$ such that $j\neq i$. Then, define positive local costs $r_{k}(u,v)$ and $r_{k}(v,w)$ such that $r_{k}(u,v)+r_{k}(v,w)=r_{k}(u,w)$ for all $k\in[n]$.
1.3 Shortest path games and Nash equilibria
The considered model can be naturally interpreted in terms of game theory. Digraph $\Gamma=(V,E)$ models an $n$-person positional game: its vertex $v\in V$ is a position and a directed edge $e=(u,w)\in E$ is a legal move from $u$ to $w$. Positions $V_{T}$ are terminals, while in $V_{i}$ player $i\in[n]$ chooses a move $(v,w)\in E$, after this move the play comes from $v$ to $w$. The game starts in a fixed position $s$ and terminates in $V_{T}$.
For each $i\in[n]$, a mapping $r_{i}:E\rightarrow\mathbb{R}$ assigns a real number $r_{i}(e)$ to each move $e\in E$. This number $r_{i}(e)$ is called the local cost and it is interpreted as a cost of move $e$ for player $i\in[n]$.
A directed path $p$ from $s$ to $V_{T}$ is called a play. Its effective cost $R_{i}(p)$ for each player $i\in[n]$ is additive: $R_{i}(p)=\sum_{e\in p}r_{i}(e)$. All $n$ players are minimizers, that is, they want to reduce their effective costs.
This model defines an $n$-person SP game. It is called finite if its digraph $\Gamma$ is finite and positive if its local costs are positive: $r_{i}(e)>0$ for all players $i\in[n]$ and moves $e\in E$.
A strategy $\sigma_{i}\in S_{i}$ of a player $i\in[n]$ is a mapping $\sigma_{i}:V_{i}\rightarrow V$ that assigns to each position $v\in V_{i}$ a move $(v,\sigma_{i}(v))\in E$ from $v$. An $n$-tuple $\sigma=(\sigma_{i}\mid i\in[n])$ is called a strategy profile or a situation. Every situation $\sigma$ uniquely defines a play $p(\sigma)$ that begins in $s$ and ends at $V_{T}$; each move $(v,w)$ of $p(\sigma)$ is determined by the corresponding strategy $w=\sigma_{i}(v)$ if $v\in V_{i}$.
A situation $\sigma=(\sigma_{1},\dots,\sigma_{n})$ is called a Nash Equilibrium (NE) if no player $i\in[n]$ can improve his/her result $R_{i}(p(\sigma))$ by replacing his/her strategy $\sigma_{i}$ by some other strategy $\sigma^{\prime}_{i}$ provided $n-1$ remaining players keep their strategies unchanged, that is, $R(p(\sigma))\leq R(p(\sigma^{\prime}))$ provided $\sigma^{\prime}$ may differ from $\sigma$ only in the entry $i$. In other words, for each player $i\in[n]$ his/her strategy $\sigma_{i}$ is the best response to the remaining $n-1$ strategies of $\sigma$. This means stability provided only one player (not a coalition) may try to deviate from the considered situation $\sigma$. This is the simplest and most popular concept of solution in the non-cooperative game theory; it was introduced in 1950 by John Nash [13, 14].
A NE $\sigma$ is called terminal if the corresponding play $p(\sigma)$ is a path from $s$ to $V_{T}$ and cyclic if $p(\sigma)$ starts in $s$ and does not come to $V_{T}$. In this case it is a “lasso” that starts in $s$, proceed with an initial path, which may be empty, and cycles as soon as it revisits a position.  [^3] In this case it will repeat the obtained directed cycle infinitely, because we restrict the players by they stationary strategies, which repeat a move whenever a position is revisited.
For a positive SP games, all cyclic plays are of infinite cost for each player. We treat them as equivalent and denote by the same symbol $\infty$. Thus, every terminal play is cheaper than any cyclic one for each player.
By definition, if $p\in P_{1}\cap P_{2}$ then there exists a situation $\sigma=(\sigma_{1},\sigma_{2})$ such that $p=p_{2}(\sigma_{1})=p_{1}(\sigma_{2})$, where $p_{3-i}(\sigma_{i})\in P_{3-i}(\sigma_{i})$, for $i=1,2$. Hence, $\sigma_{1}$ is a best response to $\sigma_{2}$ and vice versa, that is, $\sigma$ is a terminal NE. It is easily seen that we can reformulate Theorem 1 and Conjecture 2 in the game theoretic terms as follows.
Theorem 3
Every finite positive two-person SP game has a NE. Furthermore, such a NE can be computed in polynomial time.
Note, however, that this NE may be cyclic. Furthermore, a cyclic NE exists in a two-person SP game whenever each of the two players has a strategy cutting $V_{T}$ from $s$. Indeed, if each player applies such a strategy then $\infty$ is the only possible outcome and we obtain an NE, although $\infty$ is of infinite cost for both players. We conjecture that a terminal NE exists too provided $\Gamma$ has a directed path from $s$ to $V_{T}$ (an $(s,t)$-path if the reduction (j) is applied).
Conjecture 4
Every finite positive two-person SP game has a terminal NE provided it has at least one terminal play.
To prove this conjecture we have to assume that each player can cut $V_{T}$ from $s$ and construct a terminal NE; see the last section for more details.
Recall the “wlog simplifications” (jv) and (v) of Subsection 1.2. Obviously, (jv) and (v) may only destroy a NE but cannot create one. Since we are studying conditions that imply the existence of an NE, we can apply (jv) and (v) wlog.
The existence result in Theorem 3 can be extended to infinite graphs with finite out-degrees.
Theorem 5
Every positive two-person SP game where every vertex has finite out-degree has a NE.
1.4 Theorem 5 is almost precise
The two properties of an SP game required in Theorem 5 (positivity and being defined on two players) are essential for existence of an NE. A finite positive three-person NE-free SP game was constructed in [9, Figure 2 and Table 3].
A two-person finite NE-free SP game having some negative local costs was given in [3]. Moreover, such an example was obtained also for a SP game with non-negative local costs, yet, this SP game has a directed cycles whose all local costs for one of two players are 0 [3, Figures 4 and 5].  [^4] According to [5], an $n$-person SP game can be replaced by a nonnegative one, by applying a potential transformation, whenever all its directed cycles have nonnegative total cost for all players.
Finally, note that even a one-person positive but infinite SP game may be terminal NE-free (that is, Conjecture 4 does not hold for infinite graphs). Let $\Gamma$ be a caterpillar, with Local costs on the main path: $1,\frac{1}{4},\dots,\frac{1}{4^{k}},\dots$, and the local cost of the termination after $k=0,1,2,\ldots$ moves along the main path are $2,\frac{1}{2},\dots,\frac{2}{4^{k}},\dots$. The efficient cost of the play in $\Gamma$ that terminates after $k$ steps is $1+\frac{2}{3}(\frac{1}{2}+4^{-k})$. It is monotone decreasing with $k$. Hence, the obtained one-person game has no shortest play and, thus, no terminal NE.  [^5] Recall that wlog we can merge all terminal positions $V_{T}$ of the caterpillar $\Gamma$ getting a unique terminal $t$.
1.5 SP games are stochastic games with perfect information, positive local costs, and 1-total effective cost
The SP games can be viewed as a special case of the so-called deterministic stochastic games with perfect information introduced in 1957 by Gillette [6]. In [6], the limit mean effective cost was considered. Then total effective cost was introduced in [15]; see also [2] and [9, Section 5]. Then, in [1] a more general $k$-total effective cost was introduced: the limit mean and total effective costs correspond to $k=0$ and $k=1$, respectively. In all cases the local costs are arbitrary (not necessarily positive) real numbers. The existence of a NE (a saddle point) for the two-person zero-sum case is a classical result of the stochastic game theory. It was shown for $k=0$ in [6, 12], in [15, 2] for $k=1$, and in [1] for arbitrary $k$. However, this result cannot be extended to the two-person non zero-sum case. An example for $k=0$ was given in [7]; then it was extended to all $k$ in [1].
Interestingly, the SP games considered in the present paper correspond to the case $k=1$. Yet, we also assume that the local costs are positive. This assumption essentially simplifies general formulas for the effective 1-total cost (see [4, 1] and also [9, Section 5]) and, in accordance with Theorem 3 implies the existence of an NE in pure stationary strategies.
2 Proof of Theorem 3
Recall that $V_{i}$ is the set of positions controlled by player $i\in I=\{1,2\}$, $s\in V_{1}\cup V_{2}$ is a fixed initial position, and $t\not\in V_{1}\cup V_{2}$ is a terminal position. Let us denote by $E_{1}$, $E_{2}$ the sets of possible moves of the players 1 and 2, respectively. Thus, $E_{1}\cup E_{2}=E$ and, by assumption (v), $E_{1}\subseteq V_{1}\times(V_{2}\cup\{t\})$ and $E_{2}\subseteq V_{2}\times(V_{1}\cup\{t\})$. Let $r_{i}:E\mapsto\mathbb{R}$ be the length function of the moves for player $i\in I$, and recall the following main assumptions:
  • (A1)
    For all positions $u\in V_{1}\cup V_{2}$ there exists at least one $(u,t)$-path (by assumption (jjj)).
  • (A2)
    The $r_{i}$-length of any directed cycle is positive for both players $i\in I$.
A strategy for player $i\in I$ is a mapping $\sigma_{i}:V_{i}\to V_{3-i}$ such that for all $u\in V_{i}$ we have $(u,\sigma_{i}(u))\in E_{i}$. A situation is a pair $(\sigma_{1},\sigma_{2})$ of strategies of the players. In every situation there is a unique walk starting at $s$, called the play of the situation and denoted by $P(\sigma_{1},\sigma_{2})$. This play may be a finite path terminating in $t$, in which case the cost of player $i\in I$ is the $r_{i}$-length of this path, or it maybe infinite (that is terminating in a directed cycle), in which case the cost for both players are +$\infty$+ ∞, due to assumption (A2).
Theorem 6
Assume assumptions (A1) and (A2) and that the length functions satisfy the following conditions for all $i\in I$:
  • (L1)
    For all $u\in V_{i}$ we have $0=\min\{r_{i}(u,v)\mid(u,v)\in E_{i}\}$.
  • (L2)
    For all $u\in V_{i}$ there exists $(u,v)\in E_{i}$ s.t. $r_{i}(u,v)=0$ and $r_{3-i}(u,v)\leq 0$.
  • (L3)
    For all $u\in V_{i}$ there exists $(u,v)\in E_{i}$ s.t. $r_{3-i}(u,v)=0$.
Then $\Gamma$ has a terminal NE.
Proof: First, for all $i\in I$ and for all $u\in V_{i}$ we choose an outgoing edge $(u,v)\in E_{i}$ such that $r_{i}(u,v)=0$ and $r_{3-i}(u,v)\leq 0$. By Condition (L2) we can choose such edges. Denote by $H$ the subgraph formed by these edges. Consequently, $H$ does not contain a cycle, since by assumption (A2) all cycles have positive $r_{i}$-length (for both $i\in I$). Therefore, there exists a directed $(s,t)$-path in $H$. Let us denote this path by $p$ and its vertex set by $V(p)$.
Next, we define strategies for the players, as follows. For $i\in I$ and $u\in V_{i}\cap V(p)$ we define $\sigma_{i}(u)=v$ where $(u,v)$ is the edge of $p$ leaving $u$. For $u\in V_{i}\setminus V(p)$ we define $\sigma_{i}(u)=v$ where $r_{3-i}(u,v)=0$. By condition (L3) we can choose such a $v$.
Finally, we claim that the situation $(\sigma_{1},\sigma_{2})$ forms a terminal NE in $\Gamma$. Note that $p$ is the play of this situation. Since the role of the players is symmetric, assume for a contradiction that player $1$ can deviate and improve. Let us say that player $1$ deviates from $p$ and creates a new play $p^{\prime}$ such that the $r_{1}$-length of $p^{\prime}$ is strictly smaller than the $r_{1}$-length of $p$. Since both $p$ and $p^{\prime}$ are simple paths, every edge can appear at most once in each of these paths.
By our definition of $p$, the $r_{1}$-length of any edge in $p$ is nonpositive, and hence the total $r_{1}$-length of the edges of $p$ that are not edges of $p^{\prime}$ is nonpositive.
On the other hand, we have $r_{1}(x,y)\geq 0$ for all $x\in V_{1}$ and $(x,y)\in E_{1}$ by condition (L1), and have $r_{1}(x,\sigma_{2}(x))=0$ for all $x\in V_{2}\setminus p$ by our definition of $\sigma_{2}$. Thus the total $r_{1}$-length of edges of $p^{\prime}$ that are not edges of $p$ is nonnegative, contradicting our assumption that $p^{\prime}$ is a strict improvement over $p$. $\Box$
Given a length function $r:E_{1}\cup E_{2}\mapsto\mathbb{R}$ and a potential function $\phi:V_{1}\cup V_{2}\cup\{t\}\mapsto\mathbb{R}$, the potential transformation of $r$ is defined by
$\bar{r}(u,v)=r(u,v)+\phi(v)-\phi(u)~{}\text{ for all }~{}(u,v)\in E_{1}\cup E_{2}.$
It is important to note that such potential transformation does not change the $r$-costs of the cycles, or the relative order of $(s,t)$-paths with respect to the total $r$-cost. In particular, applying such transformations to the costs of the two players in an SP game does not affect the existence of a NE in the game.
We say that player $i\in I$ can block, if it has a strategy $\sigma_{i}$ such that in the subgraph $\{(u,\sigma_{i}(u))\mid u\in V_{i}\}\cup E_{3-i}$ there is a nonempty set $B\subseteq(V_{1}\cup V_{2})\setminus\{s\}$ with the property that no $(u,t)$-path exists for all $u\in B$. We say that player $i\in I$ can force +$\infty$+ ∞, if it has a strategy $\sigma_{i}$ such that in the subgraph $\{(u,\sigma_{i}(u))\mid u\in V_{i}\}\cup E_{3-i}$ there is no $(s,t)$-path. Note that a player that cannot force +$\infty$+ ∞ still may be able to block.
The next theorem is the simplest case when we can show the existence of a NE, under the strong condition that none of the players can block.
Theorem 7
Assume (A1), (A2), and that none of the players can block. Then there exist (polynomially computable) potential functions $\phi_{i}$, $i\in I$ such that the potential transformations
$\bar{r}_{i}(u,v)=r_{i}(u,v)+\phi_{i}(v)-\phi_{i}(u)~{}\text{ for all }~{}(u,v)\in E_{1}\cup E_{2}.$
satisfy conditions (L1), (L2), and (L3) for both $i\in I$.
Proof: By assumption, for both $i\in I$ and for all strategies of player $3-i$, the $r_{i}$-shortest paths have finite length for all positions. For a position $u\in V_{1}\cup V_{2}\cup\{t\}$ let us denote by $\phi_{i}(u)$ the length of the longest such $r_{i}$-shortest path, where we maximize over all strategies of player $3-i$. In fact (see [11, 10]), for each $i\in I$ there exists a single tree rooted at $t$ that represents all such shortest-longest paths, and it can be computed in polynomial time. Note that $\phi_{i}$ satisfies the following relation for all positions:
$\phi_{i}(u)=\begin{cases}\min_{(u,v)\in E_{i}}r_{i}(u,v)+\phi_{i}(v)\text{ if }u\in V_{i},\\ \max_{(u,v)\in E_{3-i}}r_{i}(u,v)+\phi_{i}(v)\text{ if }u\in V_{3-i}.\end{cases}$
It follows that for the transformed lengths we have for both $i\in I$ that
$\begin{array}[]{rl}0&=\min\bar{r}_{i}(u,v)\text{ for }u\in V_{i},\\ 0&=\max\bar{r}_{i}(u,v)\text{ for }u\in V_{3-i}.\end{array}$ (1)
Finally we note that (L1), (L2) and (L3) follow from (1). $\Box$
Theorem 8
Assume (A2) and that one of the payers cannot force +$\infty$+ ∞. Then $\Gamma$ has a terminal NE, which can be computed in polynomial time.
Proof: Wlog, we may assume that player $2$ cannot force +$\infty$+ ∞.
First we consider the $r_{1}$-shortest-longest paths, where we minimize over player $1$’s strategies, and maximize over strategies of player $2$. Let us denote by $\phi_{1}(u)$ the length of such an $r_{1}$-shortest-longest path from position $u\in V_{1}\cup V_{2}\cup\{t\}$. By [10] we can determine in polynomial time a directed tree rooted at $t$ that represents $r_{1}$-shortest-longest paths for every position for which $\phi_{1}$ is finite. Let us denote by
$B=\{u\in V_{1}\cup V_{2}\mid\phi_{1}(u)=+\infty\}$
the subset of the positions where $\phi_{1}$ is not finite, and let $U=(V_{1}\cup V_{2}\cup\{t\})\setminus B$. Note that the set $B$ can also be computed in polynomial time (see [10]). We denote by $E_{1}^{\prime}\subseteq E_{1}$ and $E_{2}^{\prime}\subseteq E_{2}$ the edge sets in the subgraph, induced by $U$.
Then we have
$\phi_{1}(u)=\begin{cases}\displaystyle\min_{v:(u,v)\in E_{1}^{\prime}}r_{1}(u,v)+\phi_{1}(v)\text{ if }u\in V_{1}\setminus B,\\ \displaystyle\max_{v:(u,v)\in E_{2}^{\prime}}r_{1}(u,v)+\phi_{1}(v)\text{ if }u\in V_{2}\setminus B.\end{cases}$ (2)
Note that by our assumptions, we have $s\in U$. Furthermore, player $1$ cannot leave the set $B$, while player $2$ cannot enter it (from $U$). Thus, we have the following properties:
$\begin{array}[]{l}\forall u\in V_{1}\cap B,~{}(u,v)\in E_{1}\text{ we have }v\in B,\\ \nexists u\in V_{2}\cap U,~{}(u,v)\in E_{2}\text{ with }v\in B,\\ \forall u\in V_{2}\cap B~{}\exists(u,v)\in E_{2}\text{ with }v\in B.\end{array}$ (3)
We then use $\phi_{1}$ and apply potential transformation to the $r_{1}$ length function in the subgraph induced by $U$. Then for the transformed lengths we have by (2) that
00 0 $=\displaystyle\min_{(u,v)\in E_{1}^{\prime}}\bar{r}_{1}(u,v)\text{ for }u\in V_{1}\setminus B,$ (4a)
00 0 $=\displaystyle\max_{(u,v)\in E_{2}^{\prime}}\bar{r}_{1}(u,v)\text{ for }u\in V_{2}\setminus B.$ (4b)
Note that we have no precise information about $\bar{r}_{1}(u,v)$ for $(u,v)\notin E_{1}^{\prime}\cup E_{2}^{\prime}$. Nevertheless, by equations (4) we have for all $u\in(V_{1}\cup V_{2})\setminus B$ a move $(u,v)\in E_{1}^{\prime}\cup E_{2}^{\prime}$ such that $\bar{r}_{1}(u,v)=0$. For $u\in V_{1}$, let us define
$\sigma_{1}(u)=\begin{cases}v&\text{ for }u\in V_{1}\setminus B\text{ such that}(u,v)\in E_{1}^{\prime},~{}\bar{r}_{1}(u,v)=0,\\ v&\text{ for }u\in B\text{ arbitrarily such that }(u,v)\in E_{1}.\end{cases}$ (5)
By (3) and (4a) strategy $\sigma_{1}$ is well-defined. We introduce the notation
$H=\{(u,\sigma_{1}(u))\mid u\in V_{1}\setminus B\}$ (6)
for the set of $\sigma_{1}$-moves of player $1$ within $U$.
The subgraph $(U,H\cup E_{2}^{\prime})$ now contains the spanning $r_{1}$-shortest-longest paths we constructed within $E_{1}^{\prime}\cup E_{2}^{\prime}$. Thus the $r_{2}$ shortest $(u,t)$-path in the subgraph $(U,H\cup E_{2}^{\prime})$ is finite for each $u\in U$.
Let us denote by $p$ an $r_{2}$-shortest $(s,t)$-path, and for $u\in V_{2}$, define
$\sigma_{2}(u)=\begin{cases}v&\text{ for }u\in V(p)\cap V_{2}\text{ such that }(u,v)\text{ is an edge of }p,\\ v&\text{ for }u\in V_{2}\setminus(B\cup V(p))\text{ such that }\bar{r}_{1}(u,v)=0,\\ v&\text{ for }u\in V_{2}\cap B\text{ such that }v\in B,~{}(u,v)\in E_{2}.\end{cases}$
By (4b) we have a $v$ for all $u\in V_{2}\setminus V(p)$ such that $\bar{r}_{1}(u,v)=0$. By (3) for every $u\in V_{2}\cap B$ we have a $(u,v)\in E_{2}$ such that $v\in B$. Thus, $\sigma_{2}$ is a well-defined strategy of player $2$.
We claim that the situation $(\sigma_{1},\sigma_{2})$ is a terminal NE. Note that $p$ is the play of this situation, and hence it is terminal.
Let us first observe that, by (6), player $2$ could improve on $p$ only if there exists an $(s,t)$-path within the subgraph $(U,H\cup E_{2}^{\prime})$, which has a shorter $r_{2}$-length than $p$. This is however impossible, since $p$ was chosen as an $r_{2}$-shortest $(s,t)$-path. Furthermore, by (3) player $2$ cannot enter $B$. Thus player $2$ cannot change $\sigma_{2}$ and improve strictly.
As for player $1$, let us note first that for all $u\in V(p)\cap V_{1}$ we have $\bar{r}_{1}(u,\sigma_{1}(u))=0$ and for all $v\in V_{2}\cap V(p)$ we have $\bar{r}_{1}(v,\sigma_{2}(v))\leq 0$ by our definitions of $\sigma_{1}$ and $\sigma_{2}$, and by (4b). Thus for player $1$ every move along $p$ has a nonpositive $\bar{r}_{1}$-length.
For his moves outside $p$, we have $\bar{r}_{1}(u,\sigma_{2}(u))=0$ for all $u\in V_{2}\setminus(B\cup V(p))$ by our definition of $\sigma_{2}$, and for all moves $(u,v)\in E_{1}^{\prime}$, $u\in V_{1}\cap U$ we have $\bar{r}_{1}(u,v)\geq 0$ by (4a). Thus all player $1$ can do is to replace some nonpositive moves along $p$ by some nonnegative moves outside $p$, and hence he cannot improve either. This proves our claim that $(\sigma_{1},\sigma_{2})$ is a NE, completing the proof of the theorem. $\Box$
As mentioned earlier, when both players can force +$\infty$+ ∞, then the pair of forcing strategies constitutes a (cyclic) NE. As can be seen from the proof of Theorem 8, such +$\infty$+ ∞-forcing strategies for the two players can be found in polynomial time by computing the $r_{i}$-shortest-longest paths, and checking if the lengths obtained satisfy $\phi_{i}(s)=+\infty$, for $i\in I$. This concludes the proof of Theorem 3.
Remark 1
Let us add that we could have used the lengths of the shortest $r_{2}$-paths in the second part of the above proof as a potential function, and transform with it the $r_{2}$-lengths of the edges in $E_{2}^{\prime}$. It could then be shown that in this situation properties (L1), (L2), and (L3) hold in the subgraph induced by $U$. Thus, there exists a terminal NE within this subgraph by Theorem 6. Using properties (3) then we could have argued that this terminal NE can be extended to a NE in $\Gamma$.
3 Proof of Theorem 5
Due to compactness arguments, Theorem 1 holds also in the case of infinite digraphs such that all vertices have a finite out-degree. In this case infinite plays may be infinite paths with finite effective costs.
Let $\Gamma=(V,E)$ be an infinite digraph, $s$ be the initial vertex. We assume that $V$ is countable and fix an enumeration of vertices: $v_{0},v_{1},\dots$, where $v_{0}=s$. For each $n$ define a finite part of the game $\Gamma=(V^{n},E^{n})$ to be a finite digraph on the set
$V^{n}=\bigcup_{i=0}^{n-1}\big{(}\{v_{i}\}\cup N^{+}(v_{i})\big{)},\quad\text{where}\ N^{+}(v)=\{u:(v,u)\in E\}.$
having the edge set
$E^{N}=\{(v_{j},v_{r}):0\leq j<n,(v_{j},v_{r})\in E\}.$
Players control the same vertices in the set $\{v_{0},\dots,v_{n-1}\}$ as in the infinite game $\Gamma$ and vertices $v_{r}$, $r\geq n$, are terminals in $\Gamma^{n}$. Local costs are inherited from $\Gamma$.
By Theorem 1, there exist NE for each $\Gamma^{n}$. Fix NE strategy profiles $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$ for each $\Gamma^{n}$. Construct a pair $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$ of strategies for infinite game such that for all $m$ and for infinitely many values of $n$ the restriction of $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$ on the game $\Gamma^{m}$ coincides with the restriction of $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$ on the game $\Gamma^{m}$. It can be done inductively. For each $m$ there are finitely many strategy profiles for $\Gamma^{m}$. Hence it is possible to choose $(\sigma^{m}_{1},\sigma^{m}_{2})$ such that $(\sigma^{m}_{1},\sigma^{m}_{2})$ is the restriction of $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$ for infinitely many $n$. Once $(\sigma^{m}_{1},\sigma^{m}_{2})$ is chosen, we continue for $m^{\prime}>m$ choosing among strategy profiles whose restrictions on $\Gamma^{m}$ coincide with $(\sigma^{m}_{1},\sigma^{m}_{2})$.
We claim that $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$ is NE in $\Gamma$. Suppose for the sake of contradiction that player $i$ can deviate and improve. For a path $p$ in the game $\Gamma$ we denote by $p[m]$ the initial part of $p$ inside the game $\Gamma^{m}$ (it finishes when $p$ enters first time a vertex $v_{j}$, $j\geq m$). Let $p$ be the deviated path; $\tilde{p}$ is the path generated by $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$; and $\tilde{p}^{n}$ be the path generated by $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$.
By construction of $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$, for any $t$, the path $\tilde{p}[t]$ is the initial part of the path $\tilde{p}^{n}$ for infinitely many $n$. Thus $R_{i}(\tilde{p}^{n})\geq R_{i}(\tilde{p}[t])$ for infinitely many $n$. Note that $R_{i}(p[n])\geq R_{i}(\tilde{p}^{n})$, since $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$ is NE in the game $\Gamma^{n}$. Therefore, for all $t$ and infinitely many $n$, $R_{i}(p[n])\geq R_{i}(\tilde{p}[t])$. Since $R_{i}(p)=\lim_{n\to\infty}R_{i}[n]$ and $R_{i}(\tilde{p})=\lim_{t\to\infty}R_{i}(\tilde{p}[t])$, we get $R_{i}(p)\geq R_{i}(\tilde{p})$, which contradicts to the assumption that player $i$ improves.
Acknowledgements
The third and forth authors were working within the framework of the HSE University Basic Research Program.