### v4

# Two-person Positive Shortest Path Games Have Nash Equilibria in Pure Stationary Strategies

Abstract
We prove that every finite two-person shortest path game, where the local cost of every move is positive for each player, has a Nash equilibrium (NE) in pure stationary strategies, which can be computed in polynomial time. We also extend the existence result to infinite graphs with finite out-degrees. Moreover, our proof gives that a terminal NE (in which the play is a path from the initial position to a terminal) exists provided at least one of the two players can guarantee reaching a terminal. If none of the players can do it, in other words, if each of the two players has a strategy that separates all terminals from the initial position $s$, then, obviously, a cyclic NE exists, although its cost is infinite for both players, since we restrict ourselves to positive games. We conjecture that a terminal NE exists too, provided there exists a directed path from $s$ to a terminal. However, this is open.
We extend our result to short paths interdiction games, where at each vertex, we allow one player to block some of the arcs and the other player to choose one of the non-blocked arcs. Assuming that blocking sets are chosen from an independence system given by an oracle, we give an algorithm for computing a NE in time $O(|E|(\log|V|+\tau))$, where $V$ is the set of vertices, $E$ is the set of arcs, and $\tau$ is the maximum time taken by the oracle on any input.
Keywords: two-person shortest path games, Nash equilibrium in pure stationary strategies, short paths interdiction, independence system.
1 Introduction
1.1 Shortest Path Games and Nash Equilibria
Let $G=(V,E)$ be a finite directed graph (digraph) that may have loops, multiple arcs, and directed cycles. A fixed vertex $s\in V$ is called initial. Furthermore, $V$ is partitioned into $3$ subsets $V=V_{1}\cup V_{2}\cup V_{T}$ such that $E$ contains no arc $(u,v)$ leaving $u$ if and only if $u\in V_{T}$. In the latter case $v$ is called terminal or a terminal vertex.
For each $i\in\{1,2\}$, a mapping $r_{i}:E\rightarrow\mathbb{R}$ assigns a real number $r_{i}(e)$ to each arc $e\in E$, called the $i$th local cost (or length) of $e$. Given $i\in\{1,2\}$ and a directed path $p$ from $s$ to a terminal $v\in V_{T}$, we assign the real number $R_{i}(p)=\sum_{e\in p}r_{i}(e)$ and call it the $r_{i}$-effective cost (distance, or length) of $p$.
The considered model can be naturally interpreted in terms of game theory. Digraph $G=(V=(V_{1}\cup V_{2}),E)$ models a two-person positional game $\Gamma=(G,s,V_{T},r_{1},r_{2})$: its vertex $v\in V_{i}$ is a position controlled by player $i\in\{1,2\}$, and an arc $e=(u,v)\in E$ is a legal move from $u$ to $v$. The game starts in a fixed (initial) position $s$; when the control is at vertex $u\in V_{i}$, player $i\in\{1,2\}$ chooses a move $(u,v)\in E$, after which the control moves to $v$. The game terminates when a vertex $t\in V_{T}$ is reached. The $i$th local cost $r_{i}(e)$ is interpreted as a cost of move $e$ for player $i\in\{1,2\}$.
A play $p$ in this game is defined to be either a directed path from $s$ to some vertex $t\in V_{T}$ (called $(s,t)$-path), or as an “$s$-lasso”, that is, a (possibly empty) directed path starting in $s$ and ending in a directed cycle. If the play is a path, then its effective cost $R_{i}(p)$ for each player $i\in[n]$ is assumed to be additive: $R_{i}(p)=\sum_{e\in p}r_{i}(e)$; if $p$ is an $s$-lasso, then $R_{i}(p)=+\infty$, 00, or $\infty$- ∞, depending on whether $R_{i}(p)>0$, $R_{i}(p)=0$ or $R_{i}(p)<0$, respectively. We assume that both players are minimizers, that is, they want to reduce their effective costs.
We call such a game a two-person shortest path (SP) game. It is said to be finite if its digraph $G$ is finite and positive if its local costs are positive [^1] Equivalently, all directed cycles have positive total costs for each player [8].: $r_{i}(e)>0$ for each player $i\in\{1,2\}$ and each move $e\in E$.
A strategy $\sigma_{i}$ of a player $i\in\{1,2\}$ is a mapping $\sigma_{i}:V_{i}\rightarrow V$ that assigns to each position $v\in V_{i}$ a move $(v,\sigma_{i}(v))\in E$ from $v$. We denote by $S_{i}$ the set of strategies of player $i\in\{1,2\}$. A pair $\sigma=(\sigma_{1},\sigma_{2})$ is called a strategy profile or a situation. Every situation $\sigma$ uniquely defines a play $p(\sigma)$; each move $(u,v)$ of $p(\sigma)$ is determined by the corresponding strategy $v=\sigma_{i}(u)$ if $u\in V_{i}$.
A situation $\sigma=(\sigma_{1},\sigma_{2})$ is called a Nash Equilibrium (NE) if no player $i\in\{1,2\}$ can improve his/her effective cost $R_{i}(p(\sigma))$ by replacing his/her strategy $\sigma_{i}$ by some other strategy $\sigma^{\prime}_{i}$, provided the other player keeps his/her strategy unchanged, that is, for all $i\in\{1,2\}$, $R_{i}(p(\sigma)\leq R_{i}(p(\sigma^{\prime}))$, provided $\sigma^{\prime}$ may differ from $\sigma$ only in the $i$th entry. In other words, for each player $i\in\{1,2\}$ his/her strategy $\sigma_{i}$ is the best response to the other player’s strategy $\sigma_{3-i}$. This means stability provided only one player (not a coalition) may try to deviate from the considered situation $\sigma$. This is the simplest and most popular concept of solution in the non-cooperative game theory; it was introduced in 1950 by John Nash [20, 21].
A NE $\sigma$ in an SP game is called terminal if the corresponding play $p(\sigma)$ is a path from $s$ to $V_{T}$ and cyclic if $p(\sigma)$ starts in $s$ and does not come to $V_{T}$. In this case it is an $s$-lasso that starts in $s$, proceeds with an initial path, which may be empty, and cycles as soon as it revisits a position.  [^2] In this case it will repeat the obtained directed cycle infinitely, because we restrict the players by their stationary strategies, which repeat a move whenever a position is revisited.
For a positive SP game, all cyclic plays are of infinite cost for each player. We treat them as equivalent and denote any of them by the same symbol $\overline{\infty}$. Thus, every terminal play is cheaper than any cyclic one for each player. For $i\in\{1,2\}$, we say that player $i$ can cut $V_{T}$ from $s$, if he/she has a strategy $\sigma_{i}$ such that the graph obtained from $G$ by fixing $\sigma_{i}$ (and deleting all arcs not selected by $\sigma_{i}$ at $v\in V_{i}$) does not contain a directed path from $s$ to any of the terminals (called $(s,V_{T})$-path).
The main result of this paper is the following.
Theorem 1
Every finite positive two-person SP game has a NE. Furthermore, such a NE can be computed in polynomial time.
Note, however, that this NE may be cyclic. Furthermore, a cyclic NE exists in a two-person SP game whenever each of the two players has a strategy cutting $V_{T}$ from $s$. Indeed, if each player applies such a strategy then $\overline{\infty}$ is the only possible outcome and we obtain an NE, although $\overline{\infty}$ is of infinite cost for both players. We conjecture that a terminal NE exists too provided $G$ has a directed path from $s$ to $V_{T}$.
Conjecture 2
Every finite positive two-person SP game has a terminal NE provided it has at least one terminal play.
[…]
Finally, we observe that the existence result in Theorem 1 can be extended to infinite graphs with finite out-degrees.
Theorem 3
Every positive two-person SP game where every vertex has finite out-degree has a NE.
Variants of shortest path games have been considered extensively in the literature, see, e.g., [5, 6, 7, 22, 23, 26, 28, 29], mostly focusing on pure history dependent strategies and solving by backward induction.
1.2 Bi-shortest path Theorem and Conjecture
The game-theoretic result stated above can also be interpreted purely in terms of graph theory. As before, let $G=(V,E)$ be a digraph, $s\in V$ be a given vertex, and $V_{1}\cup V_{2}\cup V_{T}$ be a partition of $V$ into $3$ subsets such that all vertices in $V_{T}$ are terminals. For each $i\in\{1,2\}$, $r_{i}:E\rightarrow\mathbb{R}$ be a mapping that assigns a local cost (or length) $r_{i}(e)$ to each arc $e\in E$, and for a directed path $p$ from $s$ to a terminal $v\in V_{T}$, define the effective (or total) cost (or length) of $p$ to be $R_{i}(p)=\sum_{e\in p}r_{i}(e)$. Denote further by $\sigma_{i}:V_{i}\rightarrow V$ a mapping that assigns to each vertex $u\in V_{i}$ a vertex $v=\sigma_{i}(u)\in V$ such that $(u,v)\in E$ is an arc of $G$, and let $S_{i}$ be the set of such mappings.
Assume that all local costs are positive ($r_{i}(e)>0$ for all $i\in\{1,2\}$ and $e\in E$). Fix a mapping $\sigma_{1}\in S_{1}$ and for each $v\in V_{1}$, choose the arc $(v,\sigma_{1}(v))\in E$ and delete from $G$ all other arcs $(v,w)$ going from $v$. Denote the obtained reduced digraph by $G(\sigma_{1})$ and denote by $P_{2}(\sigma_{1})$ the set of all shortest paths (SP) from $s$ to (some vertex in) $V_{T}$ with respect to the local lengths $r_{2}$, or in other words, minimizing the effective cost $R_{2}(p)$ over all $(s,V_{T})$-paths $p$.
It may happen that $\sigma_{1}$ cuts $V_{T}$ from $s$, that is, digraph $G(\sigma_{1})$ has no $(s,V_{T})$-paths. In this case we set $P_{2}(\sigma_{1})=\overline{\infty}$, where $\overline{\infty}$ is a special symbol, denoting a path of effective cost +$\infty$+ ∞. Set $P_{2}=\cup_{\sigma_{1}\in S_{1}}P_{2}(\sigma_{1})$. Then, by swapping the indices $1$ and $2$ define $P_{1}$ similarly.
It is easily seen that we can reformulate Theorem 1 and Conjecture 2 in the graph theoretic terms as follows.
Theorem 4 (Bi-SP Theorem)
$P_{1}\cap P_{2}\neq\emptyset$.
In other words, sets $P_{1}$ and $P_{2}$ always have a directed path or symbol $\overline{\infty}$ in common.
Clearly Theorem 4 follows immediately from Theorem 1 if we recal the game-theoretic interpretation: if $p\in P_{1}\cap P_{2}$ then there exists a situation $\sigma=(\sigma_{1},\sigma_{2})$ such that $p=p_{2}(\sigma_{1})=p_{1}(\sigma_{2})$, where $p_{3-i}(\sigma_{i})\in P_{3-i}(\sigma_{i})$, for $i=1,2$. Hence, $\sigma_{1}$ is a best response to $\sigma_{2}$ and vice versa, that is, $\sigma$ is a terminal NE.
By definition, ¯$\overline{\infty}\in P_{i}$ for $i=1,2$ if and only if there exists a mapping $\sigma_{3-i}$ cutting $V_{T}$ from $s$, that is, the reduced digraph in $G(\sigma_{3-i})$ has no $(s,V_{T})$-path. In particular, ¯$\overline{\infty}\in P_{1}\cap P_{2}$ if and only if such cutting mappings $\sigma_{1}$ and $\sigma_{2}$ both exist. We conjecture that even in this case $P_{1}$ and $P_{2}$ have an $(s,V_{T})$-path in common provided it exists in $G$.
Conjecture 5 (Bi-SP Conjecture [3, 12, 13])
If $G$ has an $(s,V_{T})$-path then sets $P_{1}$ and $P_{2}$ have such a path in common, that is, $(P_{1}\cap P_{2})\setminus\{\overline{\infty}\}\neq\emptyset$.
Note that the case when $G$ has no $(s,V_{T})$-path is trivial. Indeed, in this case all considered sets contain only $\overline{\infty}$ and no other elements.
As we know, ¯$\overline{\infty}\not\in P_{i}$ if and only if no $\sigma_{3-i}\in S_{3-i}$ cuts $V_{T}$ from $s$. If this happens for $i=1$ or $i=2$ then the Bi-SP Theorem and Conjecture coincide. Thus, the latter remains open only if there exist both $\sigma_{1}\in S_{1}$ and $\sigma_{2}\in S_{2}$ cutting $V_{T}$ from $s$. Otherwise, the statement of Conjecture 5 follows from Theorem 4.
1.3 Simplifying Assumptions
Without loss of generality (wlog) we can assume the following.
  • (i)
    $|V_{T}|=1$. Indeed, merge all terminal vertices in $V_{T}$ and replace them by a unique terminal $t$. After this, a directed path from $s$ to $V_{T}$ will be called an $(s,t)$-path.
  • (ii)
    An $(s,t)$-path exists. Indeed, otherwise $P_{1}=P_{2}=\{\overline{\infty}\}$.
  • (iii)
    Each vertex $v\in V$ and arc $e\in E$ is reachable from $s$. Indeed, otherwise it can be just deleted.
The next two simplifying assumptions also can be made without loss of generality and, although are not necessarily used in the proofs, they may help the reader to have a better understanding of the problem.
  • (iv)
    All directed $(s,t)$-paths have pairwise distinct effective costs. Indeed, we can get rid of all ties by an arbitrary small perturbation of local costs. After this, a shortest $(s,t)$-path is unique in $G(\sigma_{i})$, that is, $|P_{i}(\sigma_{3-i})|=1$ for $i=1,2$.
  • (v)
    The induced subgraph $G[V\setminus\{t\}]$ is bipartite and for every arc $(u,w)$ in it, either $u\in V_{1},w\in V_{2}$ or vice versa. In other words, two players make moves alternately. Assume that $u,w\in V_{1}$ (resp., $u,w\in V_{2})$. Then, subdivide this arc into two replacing $(u,w)$ by $(u,v)$ and $(v,w)$, where $v\in V_{2}$ (resp., $v\in V_{1})$ and define positive local costs $r_{k}(u,v)$ and $r_{k}(v,w)$ such that $r_{k}(u,v)+r_{k}(v,w)=r_{k}(u,w)$ for both $k\in\{1,2\}$.
Obviously, (iv) and (v) may only destroy a NE but cannot create one. Since we are studying conditions that imply the existence of an NE, we can apply (iv) and (v) wlog.
1.4 Necessary Conditions for Theorem 3 and Conjecture  2
The two properties of an SP game required in Theorem 3 (positivity and being defined on two players) are essential for existence of an NE. A finite positive three-person NE-free SP game was constructed in [13, Figure 2 and Table 3].
A two-person finite NE-free SP game having some negative local costs was given in [3]. Moreover, such an example was obtained also for an SP game with non-negative local costs, yet, this SP game has directed cycles whose all local costs for one of the two players are 00 [3, Figures 4 and 5].  [^4] According to [8], an $n$-person SP game can be replaced by a nonnegative one, by applying a potential transformation, whenever all its directed cycles have nonnegative total cost for all players.
Finally, note even a one-person positive but infinite SP game may be terminal NE-free (that is, Conjecture 2 does not hold for infinite graphs). Let $G$ be a caterpillar, with local costs on the main path: $1,\frac{1}{4},\dots,\frac{1}{4^{k}},\dots$, and the local cost of the termination after $k=0,1,2,\ldots$ moves along the main path are $2,\frac{1}{2},\dots,\frac{2}{4^{k}},\dots$. The efficient cost of the play in $G$ that terminates after $k$ steps is $1+\frac{2}{3}(\frac{1}{2}+4^{-k})$. It is monotone decreasing with $k$. Hence, the obtained one-person game has no shortest play and, thus, no terminal NE.  [^5] Recall that wlog (assumption (i)) we can merge all terminal positions $V_{T}$ of the caterpillar $G$ getting a unique terminal $t$.
1.5 SP Games Are Stochastic Games with Perfect Information, Positive Local Costs, and 1-Total Effective Cost
The SP games can be viewed as a special case of the so-called deterministic stochastic games with perfect information introduced in 1957 by Gillette [9]. In [9], the limit mean effective cost was considered. Then total effective cost was introduced in [27]; see also [2] and [13, Section 5]. Then, in [1] a more general $k$-total effective cost was introduced: the limit mean and total effective costs correspond to $k=0$ and $k=1$, respectively. In all cases the local costs are arbitrary (not necessarily positive) real numbers. The existence of a NE (a saddle point) for the two-person zero-sum case is a classical result of the stochastic game theory. It was shown for $k=0$ in [9, 18], in [27, 2] for $k=1$, and in [1] for arbitrary $k$. However, this result cannot be extended to the two-person non zero-sum case. An example for $k=0$ was given in [11]; then it was extended to all $k$ in [1].
Interestingly, the SP games considered in the present paper correspond to the case $k=1$. Yet, we also assume that the local costs are positive. This assumption essentially simplifies general formulas for the effective 1-total cost (see [4, 1] and also [13, Section 5]) and, in accordance with Theorem 1 implies the existence of an NE in pure stationary strategies.
1.6 Two-Person Short Paths Interdiction Games
Given a digraph $G=(V,E)$, envisioned as a network, with a non-negative cost (length) function on its arcs $r:E\rightarrow\mathbb{R}_{+}$ and two terminals $s,t\in V$, an attacker may want to destroy all short directed paths from $s$ to $t$ in $G$ by eliminating some arcs of $A$, given some limited resource that he can use for such destruction. This problem was considered in [10, 14, 15, 17, 16, 19, 24, 30, 31, 32] under the name of the short paths interdiction (or inhibition) problem. It can be viewed as a zero-sum game. Here we extend it to the general two-person, not necessarily zero-sum, case.
In particular, we consider a generalization of the two-person SP game introduced in Section 1.1, where rather than having a partition of the vertex set $V$ into $V_{1}$ and $V_{2}$, we allow one player, say player $1$, to block some of the arcs at each vertex, and the other player to choose one of the non-blocked arcs. More formally, denote by $E(u)=\{(u,v)\in E~{}|~{}v\in V\}$ the set of arcs going out of vertex $u\in V$, and assume that at each vertex $u\in V\setminus\{t\}$, we are given an independence (blocking) system, that is, a subfamily of $E(u)$ closed under taking subsets. Let $\mathcal{I}(u)\subseteq 2^{E(u)}$ be the family of independent sets of the blocking system at $u$, and denote by $\mathcal{D}(u)$ the family of dependent sets (i.e., those that are not contained in any independent set in $\mathcal{I}(u)$). A strategy of player $1$ (resp., player $2$) is a mapping $\sigma_{1}:V\to 2^{E}$ (resp., $\sigma_{2}:V\to 2^{E}$) which assigns to $u\in V\setminus\{t\}$ an independent set $\sigma_{1}(u)\in\mathcal{I}(u)$ (resp., a dependent set $\sigma_{2}(u)\in\mathcal{D}(u)$). We assume that $E(u)\not\in\mathcal{I}(u)$, and hence $\mathcal{D}(u)\neq\emptyset$, for all $u\in V\setminus\{t\}$. As before, given fixed start and terminal vertices $s$, $t$ and a pair of positive local cost functions $r_{1},r_{2}:E\to\mathbb{R}_{+}$, of players $1$ and $2$, respectively, the objective of player $i\in\{1,2\}$ is to “agree” with player $3-i$ on a path that minimizes the $r_{i}$-distance from $s$ to $t$ in the subgraph $G(\sigma_{1},\sigma_{2})=\big{(}V,\bigcup_{v\in V\setminus\{t\}}\sigma_{2}(v)\setminus\sigma_{1}(v)\big{)}$. More precisely, for $i\in\{1,2\}$, let $P_{i}(\sigma_{1},\sigma_{2})$ be the set of $r_{i}$-shortest paths in the graph $G(\sigma_{1},\sigma_{2})$. Then the cost of player $i$ corresponding to the situation $\sigma=(\sigma_{1},\sigma_{2})$ is defined to be $r_{i}(p)$ if there is a path $p\in P_{1}(\sigma)\cap P_{2}(\sigma)$ and +$\infty$+ ∞, otherwise. The assumptions that $\sigma_{1}(v)\in\mathcal{I}(v)$ and $\sigma_{2}(v)\in\mathcal{D}(v)$, for all $v\in V$, imply that $\sigma_{2}(v)\setminus\sigma_{1}(v)\neq\emptyset$ and hence the game is well-defined. We call such a game $\Gamma=(G,s,t,\mathcal{I},r_{1},r_{2})$ an SP interdiction game. Note that we obtain an SP game as a special case if we set $\mathcal{I}(u)=\{I\subseteq E(u):~{}|I|\leq|E(u)|-1\}$ for $u\in V_{1}$ and $\mathcal{I}(u)=\{\emptyset\}$ for $u\in V_{2}$.
Remark 6
An SP game is defined above as a simultaneous-move game: at each vertex $u\in V\setminus\{t\}$, players $1$ and $2$ choose strategies $\sigma_{1}(u)\in\mathcal{I}(u)$ and $\sigma_{2}\in\mathcal{D}(u)$ simultaneously. In an alternating-move version, player $1$ chooses first strategies $\sigma_{1}(u)\in\mathcal{I}(u)$, at each $u\in V\setminus\{t\}$, then player $2$ chooses an arc $(u,v)\in E(u)\setminus\sigma_{1}(u)$, resulting in a play (which is either an $(s,t)$-path or an $s$-lasso). As we shall see later (proof of Corollary 13), the two versions are in fact equivalent, and a NE $(\sigma_{1},\sigma_{2})$ with respect to (wrt) the first definition corresponds to a NE $(\sigma_{1},\sigma_{2})$ wrt the second definition, which yields a play in the graph $G(\sigma_{1},\sigma_{2})$ defined by choosing an arbitrary path in $P_{1}(\sigma_{1},\sigma_{2})\cap P_{2}(\sigma_{1},\sigma_{2})$.
Remark 7
Even though the definition of an SP interdiction game seems to be non-symmetric with respect to the two players, it can be in fact described in a symmetric way by defining the strategies of player 1 in terms of the “dual” independence system defined by its family of dependent sets $\mathcal{I}^{*}(u)=\{E(u)\setminus I~{}|~{}I\in\mathcal{I}(u)\}$, for $u\in V$. In this case, given a pair of strategies $(\sigma_{1},\sigma_{2})$, the play is defined by an $(s,t)$-path (or an $s$-lasso) in the subgraph defined by the arcs $\bigcup_{u\in V\setminus\{t\}}\sigma_{1}(u)\cap\sigma_{2}(u)$. Equivalently, we may also define the game by the family of dependent sets $\mathcal{I}^{*}$ of player 1 and the family of independent sets $\mathcal{D}^{*}$ of player 2.
In terms of algorithms, since the number of independent sets at each vertex can be exponential, we will need to specify how the independence systems can be succinctly described. Assume that, for each vertex $u\in V\setminus\{t\}$, the family of independent sets $\mathcal{I}(u)$ is given by an independence oracle, that is, an algorithm that decides, for any given $X\subseteq E(u)$, whether $X$ is independent (equivalently, $X\in\mathcal{I}(u)$). For instance, we one can imagine a situation where the blocking of each arc $e\in E$ incurs some cost $c(e)>0$, and there is a budget of $\delta(u)$ at each vertex $u\in V\setminus\{t\}$ such that the blocking cost at $u$ cannot exceed $\delta(u).$ In such a case, the independence system at $u$ can be described by a linear inequality: $\mathcal{I}(u)=\big{\{}I\subseteq E(u)~{}|~{}\sum_{e\in I}c(e)\leq\delta(u)\big{\}}$.
Theorem 1 admits the following generalization.
Theorem 8
Every finite positive two-person SP interdiction game has a NE (in pure stationary strategies). Assuming that the independence system at each vertex is given by a polynomial time oracle, such a NE can be computed in polynomial time.
1.7 Potential Transformations
Out proofs rely heavily on the use of potential transformations defined as follows. Given a graph $G=(V,E)$, a length function $r:E\mapsto\mathbb{R}$ and a potential function $\phi:V\mapsto\mathbb{R}$, the potential transformation of $r$ is defined by
$\bar{r}(u,v)=r(u,v)+\phi(v)-\phi(u)~{}\text{ for all }~{}(u,v)\in E.$
It is important to note that such potential transformation does not change the $r$-costs of the cycles, or the relative order of $(s,t)$-paths with respect to the total $r$-cost and fixed vertices $s$ and $t$. In particular, applying such transformations to the costs of the two players in an SP game does not affect the existence of a NE in the game.
Notation.
For of a directed path $p$ in $G$ denote its vertex set by $V(p)$, and for vertices $u,w\in V(p)$, denote by $p[u,w]$ the subpath of $p$ starting at $u\in V(p)$ and ending at $w\in V(p)$.
The rest of the paper contains the proofs of Theorems 13 and 8, in Sections 23 and 4, respectively.
2 Proof of Theorem 1
Consider an SP game $\Gamma=(G=(V_{1}\cup V_{2},E),s,t,r_{1},r_{2})$. Recall that $V_{i}$ is the set of positions controlled by player $i\in\{1,2\}$, $s\in V_{1}\cup V_{2}$ is a fixed initial position, and $t\not\in V_{1}\cup V_{2}$ is a terminal position. Let us denote by $E_{1}$, $E_{2}$ the sets of possible moves of the players 1 and 2, respectively. Thus, $E_{1}\cup E_{2}=E$ and, by assumption (v), $E_{1}\subseteq V_{1}\times(V_{2}\cup\{t\})$ and $E_{2}\subseteq V_{2}\times(V_{1}\cup\{t\})$. Let $r_{i}:E\mapsto\mathbb{R}$ be the length function of the moves for player $i\in I$, and recall the following main assumptions:
  • (A1)
    For all positions $u\in V_{1}\cup V_{2}$ there exists at least one $(u,t)$-path (by assumption (iii)).
  • (A2)
    The $r_{i}$-length of any directed cycle is positive for both players $i\in\{1,2\}$.
Recall that a strategy for player $i\in\{1,2\}$ is a mapping $\sigma_{i}:V_{i}\to V_{3-i}$ such that for all $u\in V_{i}$ we have $(u,\sigma_{i}(u))\in E_{i}$. A situation is a pair $(\sigma_{1},\sigma_{2})$ of strategies of the players. In every situation there is a unique walk starting at $s$, called the play of the situation and denoted by $P(\sigma_{1},\sigma_{2})$. This play may be a finite path terminating in $t$, in which case the cost of player $i\in I$ is the $r_{i}$-length of this path, or it maybe infinite (that is terminating in a directed cycle), in which case the cost for both players are +$\infty$+ ∞, due to assumption (A2).
Lemma 9
Assume assumptions (A1) and (A2) and that the length functions satisfy the following conditions for all $i\in I$:
  • (L1)
    For all $u\in V_{i}$ we have $0=\min\{r_{i}(u,v)\mid(u,v)\in E_{i}\}$.
  • (L2)
    For all $u\in V_{i}$ there exists $(u,v)\in E_{i}$ s.t. $r_{i}(u,v)=0$ and $r_{3-i}(u,v)\leq 0$.
  • (L3)
    For all $u\in V_{i}$ there exists $(u,v)\in E_{i}$ s.t. $r_{3-i}(u,v)=0$.
Then $\Gamma$ has a terminal NE.
Proof: First, for all $i\in\{1,2\}$ and for all $u\in V_{i}$ we choose an outgoing arc $(u,v)\in E_{i}$ such that $r_{i}(u,v)=0$ and $r_{3-i}(u,v)\leq 0$. By Condition (L2) we can choose such arcs. Denote by $H$ the subgraph formed by these arcs. Consequently, $H$ does not contain a cycle, since by assumption (A2) all cycles have positive $r_{i}$-length (for both $i\in\{1,2\}$). Therefore, there exists a directed $(s,t)$-path in $H$. Let us denote this path by $p$.
Next, we define strategies for the players, as follows. For $i\in I$ and $u\in V_{i}\cap V(p)$ we define $\sigma_{i}(u)=v$ where $(u,v)$ is the arc of $p$ leaving $u$. For $u\in V_{i}\setminus V(p)$ we define $\sigma_{i}(u)=v$ where $r_{3-i}(u,v)=0$. By condition (L3) we can choose such a $v$.
Finally, we claim that the situation $(\sigma_{1},\sigma_{2})$ forms a terminal NE in $\Gamma$. Note that $p$ is the play of this situation. Since the role of the players is symmetric, assume for a contradiction that player $1$ can deviate and improve. Let us say that player $1$ deviates from $p$ and creates a new play $p^{\prime}$ such that the $r_{1}$-length of $p^{\prime}$ is strictly smaller than the $r_{1}$-length of $p$. Since both $p$ and $p^{\prime}$ are simple paths, every arc can appear at most once in each of these paths.
By our definition of $p$, the $r_{1}$-length of any arc in $p$ is nonpositive, and hence the total $r_{1}$-length of the arcs of $p$ that are not arcs of $p^{\prime}$ is nonpositive.
On the other hand, we have $r_{1}(u,V)\geq 0$ for all $u\in V_{1}$ and $(u,v)\in E_{1}$ by condition (L1), and have $r_{1}(u,\sigma_{2}(u))=0$ for all $u\in V_{2}\setminus p$ by our definition of $\sigma_{2}$. Thus the total $r_{1}$-length of arcs of $p^{\prime}$ that are not arcs of $p$ is nonnegative, contradicting our assumption that $p^{\prime}$ is a strict improvement over $p$. $\Box$
We say that player $i\in\{1,2\}$ can block, if it has a strategy $\sigma_{i}$ such that in the subgraph $\{(u,\sigma_{i}(u))\mid u\in V_{i}\}\cup E_{3-i}$ there is a nonempty set $B\subseteq(V_{1}\cup V_{2})\setminus\{s\}$ with the property that no $(u,t)$-path exists for all $u\in B$. We say that player $i\in I$ can force +$\infty$+ ∞, if it has a strategy $\sigma_{i}$ such that in the subgraph $\{(u,\sigma_{i}(u))\mid u\in V_{i}\}\cup E_{3-i}$ there is no $(s,t)$-path. Note that a player that cannot force +$\infty$+ ∞ still may be able to block.
The next result is the simplest case when we can show the existence of a NE, under the strong condition that none of the players can block.
Lemma 10
Assume (A1), (A2), and that none of the players can block. Then there exist (polynomially computable) potential functions $\phi_{i}$, $i\in\{1,2\}$ such that the potential transformations
$\bar{r}_{i}(u,v)=r_{i}(u,v)+\phi_{i}(v)-\phi_{i}(u)~{}\text{ for all }~{}(u,v)\in E_{1}\cup E_{2}.$
satisfy conditions (L1), (L2), and (L3) for both $i\in\{1,2\}$.
Proof: By assumption, for both $i\in\{1,2\}$ and for all strategies of player $3-i$, the $r_{i}$-shortest paths have finite length for all positions. For a position $u\in V_{1}\cup V_{2}\cup\{t\}$ let us denote by $\phi_{i}(u)$ the length of the longest such $r_{i}$-shortest path, where we maximize over all strategies of player $3-i$. In fact (see [17, 16]), for each $i\in\{1,2\}$ there exists a single tree rooted at $t$ that represents all such shortest-longest paths, and it can be computed in polynomial time. Note that $\phi_{i}$ satisfies the following relation for all positions:
$\phi_{i}(u)=\begin{cases}\min_{(u,v)\in E_{i}}r_{i}(u,v)+\phi_{i}(v)\text{ if }u\in V_{i},\\ \max_{(u,v)\in E_{3-i}}r_{i}(u,v)+\phi_{i}(v)\text{ if }u\in V_{3-i}.\end{cases}$
It follows that for the transformed lengths we have for both $i\in\{1,2\}$ that
$\begin{array}[]{rl}0&=\min\bar{r}_{i}(u,v)\text{ for }u\in V_{i},\\ 0&=\max\bar{r}_{i}(u,v)\text{ for }u\in V_{3-i}.\end{array}$ (1)
Finally we note that (L1), (L2) and (L3) follow from (1). $\Box$
Lemma 11
Assume (A2) and that one of the payers cannot force +$\infty$+ ∞. Then $\Gamma$ has a terminal NE, which can be computed in polynomial time.
Proof: Wlog, we may assume that player $2$ cannot force +$\infty$+ ∞.
First we consider the $r_{1}$-shortest-longest paths, where we minimize over player $1$’s strategies, and maximize over strategies of player $2$. Let us denote by $\phi_{1}(u)$ the length of such an $r_{1}$-shortest-longest path from position $u\in V_{1}\cup V_{2}\cup\{t\}$. By [17, 16] we can determine in polynomial time a directed tree rooted at $t$ that represents $r_{1}$-shortest-longest paths for every position for which $\phi_{1}$ is finite. Let us denote by
$B=\{u\in V_{1}\cup V_{2}\mid\phi_{1}(u)=+\infty\}$
the subset of the positions where $\phi_{1}$ is not finite, and let $U=(V_{1}\cup V_{2}\cup\{t\})\setminus B$. Note that the set $B$ can also be computed in polynomial time (see [16]). We denote by $E_{1}^{\prime}\subseteq E_{1}$ and $E_{2}^{\prime}\subseteq E_{2}$ the arc sets in the subgraph, induced by $U$.
Then we have
$\phi_{1}(u)=\begin{cases}\displaystyle\min_{v:(u,v)\in E_{1}^{\prime}}r_{1}(u,v)+\phi_{1}(v)\text{ if }u\in V_{1}\setminus B,\\ \displaystyle\max_{v:(u,v)\in E_{2}^{\prime}}r_{1}(u,v)+\phi_{1}(v)\text{ if }u\in V_{2}\setminus B.\end{cases}$ (2)
Note that by our assumptions, we have $s\in U$. Furthermore, player $1$ cannot leave the set $B$, while player $2$ cannot enter it (from $U$). Thus, we have the following properties:
$\begin{array}[]{l}\forall u\in V_{1}\cap B,~{}(u,v)\in E_{1}\text{ we have }v\in B,\\ \nexists u\in V_{2}\cap U,~{}(u,v)\in E_{2}\text{ with }v\in B,\\ \forall u\in V_{2}\cap B~{}\exists(u,v)\in E_{2}\text{ with }v\in B.\end{array}$ (3)
We then use $\phi_{1}$ and apply a potential transformation to the $r_{1}$ length function in the subgraph induced by $U$. Then for the transformed lengths we have by (2) that
00 0 $=\displaystyle\min_{(u,v)\in E_{1}^{\prime}}\bar{r}_{1}(u,v)\text{ for }u\in V_{1}\setminus B,$ (4a)
00 0 $=\displaystyle\max_{(u,v)\in E_{2}^{\prime}}\bar{r}_{1}(u,v)\text{ for }u\in V_{2}\setminus B.$ (4b)
Note that we have no precise information about $\bar{r}_{1}(u,v)$ for $(u,v)\notin E_{1}^{\prime}\cup E_{2}^{\prime}$. Nevertheless, by equations (4) we have for all $u\in(V_{1}\cup V_{2})\setminus B$ a move $(u,v)\in E_{1}^{\prime}\cup E_{2}^{\prime}$ such that $\bar{r}_{1}(u,v)=0$. For $u\in V_{1}$, let us define
$\sigma_{1}(u)=\begin{cases}v&\text{ for }u\in V_{1}\setminus B\text{ such that}(u,v)\in E_{1}^{\prime},~{}\bar{r}_{1}(u,v)=0,\\ v&\text{ for }u\in B\text{ arbitrarily such that }(u,v)\in E_{1}.\end{cases}$ (5)
By (3) and (4a) strategy $\sigma_{1}$ is well-defined. We introduce the notation
$H=\{(u,\sigma_{1}(u))\mid u\in V_{1}\setminus B\}$ (6)
for the set of $\sigma_{1}$-moves of player $1$ within $U$.
The subgraph $(U,H\cup E_{2}^{\prime})$ now contains the spanning $r_{1}$-shortest-longest paths we constructed within $E_{1}^{\prime}\cup E_{2}^{\prime}$. Thus the $r_{2}$-shortest $(u,t)$-path in the subgraph $(U,H\cup E_{2}^{\prime})$ is finite for each $u\in U$.
Let us denote by $p$ an $r_{2}$-shortest $(s,t)$-path, and for $u\in V_{2}$, define
$\sigma_{2}(u)=\begin{cases}v&\text{ for }u\in V(p)\cap V_{2}\text{ such that }(u,v)\text{ is an arc of }p,\\ v&\text{ for }u\in V_{2}\setminus(B\cup V(p))\text{ such that }\bar{r}_{1}(u,v)=0,\\ v&\text{ for }u\in V_{2}\cap B\text{ such that }v\in B,~{}(u,v)\in E_{2}.\end{cases}$
By (4b) we have a $v$ for all $u\in V_{2}\setminus V(p)$ such that $\bar{r}_{1}(u,v)=0$. By (3) for every $u\in V_{2}\cap B$ we have a $(u,v)\in E_{2}$ such that $v\in B$. Thus, $\sigma_{2}$ is a well-defined strategy of player $2$.
We claim that the situation $(\sigma_{1},\sigma_{2})$ is a terminal NE. Note that $p$ is the play of this situation, and hence it is terminal.
Let us first observe that, by (6), player $2$ could improve on $p$ only if there exists an $(s,t)$-path within the subgraph $(U,H\cup E_{2}^{\prime})$, which has a shorter $r_{2}$-length than $p$. This is however impossible, since $p$ was chosen as an $r_{2}$-shortest $(s,t)$-path. Furthermore, by (3) player $2$ cannot enter $B$. Thus player $2$ cannot change $\sigma_{2}$ and improve strictly.
As for player $1$, let us note first that for all $u\in V(p)\cap V_{1}$ we have $\bar{r}_{1}(u,\sigma_{1}(u))=0$ and for all $v\in V_{2}\cap V(p)$ we have $\bar{r}_{1}(v,\sigma_{2}(v))\leq 0$ by our definitions of $\sigma_{1}$ and $\sigma_{2}$, and by (4b). Thus for player $1$ every move along $p$ has a nonpositive $\bar{r}_{1}$-length.
For his/her moves outside $p$, we have $\bar{r}_{1}(u,\sigma_{2}(u))=0$ for all $u\in V_{2}\setminus(B\cup V(p))$ by our definition of $\sigma_{2}$, and for all moves $(u,v)\in E_{1}^{\prime}$, $u\in V_{1}\cap U$ we have $\bar{r}_{1}(u,v)\geq 0$ by (4a). Thus all player $1$ can do is to replace some nonpositive moves along $p$ by some nonnegative moves outside $p$, and hence he/she cannot improve either. This proves our claim that $(\sigma_{1},\sigma_{2})$ is a NE, completing the proof of the lemma. $\Box$
As mentioned earlier, when both players can force +$\infty$+ ∞, then the pair of forcing strategies constitutes a (cyclic) NE. As can be seen from the proof of Lemma 11, such +$\infty$+ ∞-forcing strategies for the two players can be found in polynomial time by computing the $r_{i}$-shortest-longest paths, and checking if the lengths obtained satisfy $\phi_{i}(s)=+\infty$, for $i\in\{1,2\}$. This concludes the proof of Theorem 1.
Remark 12
Let us add that we could have used the lengths of the shortest $r_{2}$-paths in the second part of the above proof as a potential function, and transform with it the $r_{2}$-lengths of the arcs in $E_{2}^{\prime}$. It could then be shown that in this situation properties (L1), (L2), and (L3) hold in the subgraph induced by $U$. Thus, there exists a terminal NE within this subgraph by Lemma 9. Using properties (3) then we could have argued that this terminal NE can be extended to a NE in $\Gamma$.
3 Proof of Theorem 3
Due to compactness arguments, Theorem 1 holds also in the case of infinite digraphs such that all vertices have a finite out-degree. In this case infinite plays may be infinite paths with finite effective costs.
Let $G=(V,E)$ be an infinite digraph, $s$ be the initial vertex and $\Gamma=(G,s,t,r_{1},r_{2})$ be an SP game defined on $G$. We assume that $V$ is countable and fix an enumeration of vertices: $v_{0},v_{1},\dots$, where $v_{0}=s$. For each $n$ define a finite part of the game $\Gamma^{n}=(G^{n},s,t,r_{1},r_{2})$ defined on the finite digraph $G^{n}=(V^{n},E^{n})$ with vertex set
$V^{n}=\bigcup_{i=0}^{n-1}\big{(}\{v_{i}\}\cup N^{+}(v_{i})\big{)},\quad\text{where}\ N^{+}(v)=\{u:(v,u)\in E\}.$
and having arc set
$E^{n}=\{(v_{j},v_{r}):0\leq j<n,(v_{j},v_{r})\in E\}.$
Players control the same vertices in the set $\{v_{0},\dots,v_{n-1}\}$ as in the infinite game $\Gamma$ and vertices $v_{r}$, $r\geq n$, are terminals in $\Gamma^{n}$ (and local costs are inherited from $\Gamma$).
By Theorem 4, there exists a NE for each $\Gamma^{n}$. Fix NE strategy profiles $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$ for each $\Gamma^{n}$. Construct a pair $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$ of strategies for infinite game such that for all $m$ and for infinitely many values of $n$ the restriction of $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$ on the game $\Gamma^{m}$ coincides with the restriction of $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$ on the game $\Gamma^{m}$. It can be done inductively. For each $m$ there are finitely many strategy profiles for $\Gamma^{m}$. Hence it is possible to choose $(\sigma^{m}_{1},\sigma^{m}_{2})$ such that $(\sigma^{m}_{1},\sigma^{m}_{2})$ is the restriction of $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$ for infinitely many $n$. Once $(\sigma^{m}_{1},\sigma^{m}_{2})$ is chosen, we continue for $m^{\prime}>m$ choosing among strategy profiles whose restrictions on $\Gamma^{m}$ coincide with $(\sigma^{m}_{1},\sigma^{m}_{2})$.
We claim that $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$ is a NE in $\Gamma$. Suppose for the sake of contradiction that some player $i\in\{1,2\}$ can deviate and improve. For a path $p$ in the game $\Gamma$ we denote by $p[m]$ the initial part of $p$ inside the game $\Gamma^{m}$ (it finishes when $p$ enters first time a vertex $v_{j}$, $j\geq m$). Let $p$ be the deviated path; $\tilde{p}$ is the path generated by $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$; and $\tilde{p}^{n}$ be the path generated by $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$.
By construction of $(\widetilde{\sigma}^{\infty}_{1},\widetilde{\sigma}^{\infty}_{2})$, for any $t$, the path $\tilde{p}[t]$ is the initial part of the path $\tilde{p}^{n}$ for infinitely many $n$. Thus $R_{i}(\tilde{p}^{n})\geq R_{i}(\tilde{p}[t])$ for infinitely many $n$. Note that $R_{i}(p[n])\geq R_{i}(\tilde{p}^{n})$, since $(\widetilde{\sigma}^{n}_{1},\widetilde{\sigma}^{n}_{2})$ is a NE in the game $\Gamma^{n}$. Therefore, for all $t$ and infinitely many $n$, $R_{i}(p[n])\geq R_{i}(\tilde{p}[t])$. Since $R_{i}(p)=\lim_{n\to\infty}R_{i}[n]$ and $R_{i}(\tilde{p})=\lim_{t\to\infty}R_{i}(\tilde{p}[t])$, we get $R_{i}(p)\geq R_{i}(\tilde{p})$, which contradicts to the assumption that player $i$ improves.
4 Proof of Theorem 8
4.1 Existence of NE
The following result follows from Theorem 1.
Corollary 13
Every finite positive two-person SP interdiction game has a NE (in pure stationary strategies).
Proof: We reduce any given SP interdiction game $\Gamma=(G=(V,E),\mathcal{I},r_{1},r_{2})$ to an SP game $\Gamma^{\prime}=(G^{\prime}=(V^{\prime}=V_{1}^{\prime}\cup V_{2}^{\prime},E^{\prime}),r_{1}^{\prime},r_{2}^{\prime})$ as follows (see [25] for a similar reduction in the case of zero-sum mean payoff game). For every vertex $u\in V$, and every independent set $I\in\mathcal{I}(u)$, we introduce a new vertex $u_{I}\in V^{\prime}$ and an arc $(u,u_{I})\in E^{\prime}$. We also define an arc $(u_{I},v)\in E^{\prime}$ for every pair of vertices $u,v\in V$ and independent set $I\in\mathcal{I}(u)$ such that $(u,v)\in E(u)\setminus I(u)$:
$V_{1}^{\prime}$ $=V,\ \ \ V_{2}^{\prime}=\{u_{I}:~{}u\in V,~{}I\in\mathcal{I}(u)\},\ \ \ V^{\prime}=V_{1}^{\prime}\cup V_{2}^{\prime},$
$E^{\prime}$ =$=\{(u,u_{I}):~{}v\in V\setminus\{t\},~{}I\in\mathcal{I}(v)\}\cup\{(u_{I},v):~{}u\in V\setminus\{t\},~{}I\in\mathcal{I}(u),~{}(u,v)\in E(u)\setminus I\}.$
For $i\in\{1,2\}$, we set $r_{i}^{\prime}(v,v_{I})=0$ for $v\in V\setminus\{t\},~{}I\in\mathcal{I}(v)$, and $r_{i}^{\prime}(u_{I},v)=r_{i}(u,v)$ for $(u,v)\in E(u)\setminus I,~{}I\in\mathcal{I}(u)$.
We make the following simple observations:
  • (o1)
    Any $(s,t)$-path (possibly an $s$-lasso) in $G^{\prime}$ corresponds to an $(s,t)$-path in $G$ with the same $r_{i}$-cost, for $i\in\{1,2\}$.
  • (o2)
    For every pair of strategies $(\sigma_{1},\sigma_{2})$ in $\Gamma$, we can define a pair of strategies $(\sigma_{1}^{\prime},\sigma_{2}^{\prime})$ in $\Gamma^{\prime}$ as follows: for $u\in V\setminus\{t\}$ define $\sigma_{1}(u)=u_{\sigma_{1}(u)}$, and for $I\in\mathcal{I}(u)$, define $\sigma_{2}^{\prime}(u_{I})=v$ for an arbitrary arc $(u,v)\in\sigma_{2}(u)\setminus I$.
  • (o3)
    Given a pair of strategies $(\sigma_{1}^{\prime},\sigma_{2}^{\prime})$ in $\Gamma^{\prime}$, we can define a pair of strategies $(\sigma_{1},\sigma_{2})$ in $\Gamma$ as follows: for $u\in V\setminus\{t\}$, where $\sigma_{1}^{\prime}(u)=u_{I}$, define $\sigma_{1}(u)=I$ and $\sigma_{2}(u)=\bigcup_{I^{\prime}\in\mathcal{I}(u)}\{(u,\sigma_{2}^{\prime}(u_{I^{\prime}})\}$. By construction of $G^{\prime}$, $\sigma_{2}(u)\in\mathcal{D}(u)$ since $\sigma_{2}(u)\setminus I^{\prime}\neq\emptyset$ for all $I^{\prime}\in\mathcal{I}(u)$ and $u\in V\setminus\{t\}$.
  • (o4)
[…]
  • (o5)
    If $G$ has positive local costs, then all directed cycles in $G^{\prime}$ are also positive (as all such cycles must include at least two arcs of the form $(u_{I},v)$, for some $u\in V\setminus\{t\}$, $I\in\mathcal{I}(u)$ and $v\in E(u)\setminus\mathcal{I}(u)$, which have positive costs), implying by the result in [8] that all local costs in $G^{\prime}$ can be made positive by a potential transformation.
The corollary follows from the above observations and Theorem 1. $\Box$
Note that the reduction used in the proof of Corollary 13 does not yield a polynomial-time algorithm for computing a NE in the interdiction game, since the number of independent sets at any vertex can be exponential.
4.2 A Polynomial-time Algorithm
Consider an SP interdiction game $\Gamma=(G=(V,E),s,t,\mathcal{I},r_{1},r_{2})$. Assume that, for each vertex $u\in V\setminus\{t\}$, the family of independent sets $\mathcal{I}(u)$ is given by an independence oracle and denote by $\tau$ the maximum time taken by the oracle on any set $X\subseteq E(u)$, over all $u\in V$. Any subgraph of $G$ obtained by removing a subset of arcs $I(u)\in\mathcal{I}(u)$, at every $u\in V$, will be called admissible. We recall the following result, implicitly proved in [17, 16].
Lemma 14
Given a directed graph $G=(V,E)$ with associated non-negative arc lengths $r:E\to\mathbb{R}_{+}$, and a target vertex $t$, there is an $O(|E|(\log|V|+\tau))$ algorithm (Modified Dijkstra’s) that computes a potential function $\phi:V\to\mathbb{R}$, and independent sets $\widehat{I}(u)\in\mathcal{I}(u)$, for $u\in V\setminus\{t\}$, such that
$\phi(u)$ $\geq\displaystyle r(u,v)+\phi(v)\text{ for all $(u,v)\in\widehat{I}(u)$,}$ (7a)
$\phi(u)$ $\leq r(u,v)+\phi(v),\text{ for all $(u,v)\in E(u)\setminus\widehat{I}(u)$,}$ (7b)
$\phi(u)$ $=\displaystyle r(u,v_{u})+\phi(v_{u}),\text{ for some $(u,v_{u})\in E(u)\setminus\widehat{I}(u)$,}$ (7c)
$\mathcal{I}(u)$$\not\ni\{(u,v)\in E(u)~{}|~{}\phi(u)\geq r(u,v)+\phi(v)\}.$ (7d)
Note that conditions (7b)-(7d) imply that $\phi(u)$ is the maximum shortest $r$-distance from $u$ to $t$ in any admissible subgraph of $G$. Note also that there is one admissible graph which maximizes all such distances from all vertices to $t$. For completeness, we include a proof of Lemma 14 in the appendix.
We say that player $1$ can force +$\infty$+ ∞ in an SP interdiction game, if it has a blocking strategy that prevents player $2$ from reaching the terminal starting from $s$. A terminal NE is a NE in which the costs of both players are finite.
Lemma 15
Consider an SP interdiction game $\Gamma=(G=(V,E),s,t,\mathcal{I},r_{1},r_{2})$, in which player $1$ cannot force +$\infty$+ ∞. Then $\Gamma$ has a terminal NE that can be computed in $O(|E|\log|V|+\tau|E|)$ time.
Proof: Let $\phi_{2}$ be the potential function computed in Lemma 14 with respect to the cost function $r=r_{2}$, and let $\widehat{I}(u)$, for $U\in V\setminus\{t\}$, be the obtained blocking sets. Let $B=\{u\in V~{}|~{}\phi_{2}(u)=+\infty\}$, $U=V\setminus B$, and denote by $E^{\prime}\subseteq E$ the arc set in the subgraph induced by $U$.
Note that by our assumption, we have $s\in U$. Furthermore, player $1$ has a strategy that prevents player $2$ from leaving the set $B$, but none that forces entering it (from $U$). Thus, we have the following properties:
$\begin{array}[]{l}\forall u\in B,~{}\{(u,v)\in E(u)~{}|~{}v\in U\}\in\mathcal{I}(u),\\ \forall u\in U,~{}\{(u,v)\in E(u)~{}|~{}v\in U\}\not\in\mathcal{I}(u).\end{array}$ (8)
We then use $\phi_{2}$ and apply a potential transformation to the $r_{2}$-length function in the subgraph induced by $U$. Then for the transformed lengths $\bar{r}_{2}$ we have by Lemma 14 that, for all $u\in U\setminus\{t\}$,
$\bar{r}_{2}(u,v)$ $\leq 0\displaystyle\text{ for all $(u,v)\in\widehat{I}(u)$,}$ (9a)
$\bar{r}_{2}(u,v)$ $\geq 0,\text{ for all $(u,v)\in E(u)\cap E^{\prime}\setminus\widehat{I}(u)$,}$ (9b)
$\bar{r}_{2}(u,v_{u})$ $=0\displaystyle,\text{ for some $(u,v_{u})\in E(u)\cap E^{\prime}\setminus\widehat{I}(u)$,}$ (9c)
$\mathcal{I}(u)$$\not\ni\{(u,v)\in E(u)\cap E^{\prime}~{}|~{}\bar{r}_{2}(u,v)\leq 0\}.$ (9d)
Let $E^{\prime\prime}=\{(u,v)\in E^{\prime}~{}|~{}\bar{r}_{2}(u,v)\leq 0\}$ and note that (9c) (and the absence of 00-cycles in $G$) implies that there is an $(s,t)$-path in $G^{\prime}=(U,E^{\prime\prime})$. With every $(u,v)\in E(u)\cap E^{\prime\prime}$, let us associate an independent set $I(u,v)=\{(u,v^{\prime})\in\widehat{I}(u)~{}|~{}\bar{r}_{2}(u,v^{\prime})<\bar{r}_{2}(u,v)\}$, and note by this definition that
$\bar{r}_{2}(u,v)$ $=\min_{(u,v^{\prime})\in E(u)\cap E^{\prime}\setminus I(u,v)}\bar{r}_{2}(u,v^{\prime}).$ (10)
Next, let $p$ be an $r_{1}$-shortest $(s,t)$-path in the graph $G^{\prime}$, and for $u\in V$, define
$\sigma_{1}(u)=\begin{cases}I(u,v),&\text{ for }u\in V(p)\text{ such that }(u,v)\text{ is an arc of }p,\\ \widehat{I}(u),&\text{ for }u\in U\setminus V(p),\\ \{(u,v)\in E(u)~{}|~{}v\in U\},&\text{ for }u\in B.\end{cases}$
Note that (8) implies that $\sigma_{1}$ is well-defined. For $u\in V\setminus\{t\}$, we define
$\sigma_{2}(u)=\begin{cases}E(u)\cap E^{\prime\prime},&\text{ for }u\in U,\\ D\in\mathcal{D}(u),&\text{ for }u\in B.\end{cases}$
Note by (9d) that $\sigma_{2}(u)\in\mathcal{D}(u)$. We claim that the situation $(\sigma_{1},\sigma_{2})$ is a terminal NE. Note that $p$ is an $(s,t)$-path in the graph $G(\sigma_{1},\sigma_{2})$, and hence to show that the situation $(\sigma_{1},\sigma_{2})$ is a terminal NE, it remains to prove that $p$ is an $r_{i}$-shortest path in the game $G(\sigma_{1},\sigma_{2})$.
Observe next that, by the definition of $\sigma_{2}$ and (8), player $1$ cannot force entering $B$. It follows then, by the definition of $\sigma_{1}$ and $\sigma_{2}$ that player $1$ could improve on $\sigma_{1}$ only if there exists an $(s,t)$-path within the subgraph $G^{\prime}$, which has a shorter $r_{1}$-length than $p$. This is however impossible, since $\sigma_{1}$ was chosen such that the subgrapgh $G(\sigma_{1},\sigma_{2})$ contains an $r_{1}$-shortest $(s,t)$-path. Thus player $1$ cannot change $\sigma_{1}$ and improve strictly.
As for player $2$, suppose he can improve strictly by deviating to a strategy $\sigma_{2}^{\prime}$ such that there is a path $p^{\prime}$ in the graph $G(\sigma_{1},\sigma_{2}^{\prime})$ with strictly less $r_{2}$-cost than $p$. Note that player $2$ is not interested in entering $B$ as he cannot escape from there given the choice of $\sigma_{1}$. Thus all the arcs in $p^{\prime}$ should lie inside $U$. Consider any arc $(u,v)\in p\setminus p^{\prime}$ and let $(u,v^{\prime})$ be the arc at $u$ that belongs to $p^{\prime}$. The definition of $\sigma_{1}$ and (10) imply that $\bar{r}_{2}(u,v^{\prime})\geq\bar{r}_{2}(u,v)$.
For arcs $(u,v)$ of $p^{\prime}$ such that $u\not\in V(p)$, we have $\bar{r}_{2}(u,v)\geq 0$ by (9b). Then, by the above observations and the non-positivity of the arcs on $p$,
$\bar{r}_{2}(p^{\prime}\setminus p)$ $=\sum_{(u,v)\in p^{\prime}\setminus p:~{}u\in V(p)}\bar{r}_{2}(u,v)+\sum_{(u,v)\in p^{\prime}\setminus p:~{}u\not\in V(p)}\bar{r}_{2}(u,v)$
$\geq\sum_{(u,v)\in p\setminus p^{\prime}:~{}u\in V(p^{\prime})}\bar{r}_{2}(u,v)+\sum_{(u,v)\in p\setminus p^{\prime}:~{}u\not\in V(p^{\prime})}\bar{r}_{2}(u,v)=\bar{r}_{2}(p\setminus p^{\prime}).$
Thus all player $2$ can do is to replace some moves of $p$ by some equal or costlier moves outside $p$, and hence he/she cannot improve either. This proves our claim that $(\sigma_{1},\sigma_{2})$ is a NE, completing the proof of the lemma. $\Box$
Proof of Theorem 8.
Consider an SP interdiction game $\Gamma=(G=(V,E),s,t,\mathcal{I},r_{1},r_{2})$. If both players can force +$\infty$+ ∞, then the corresponding strategies define a (cyclic) NE. Such strategies can be computed in $O(|E|\log|V|+\tau|E|)$ by applying Modified Dijkstra’s algorithm as in Lemma 14. If player 1 cannot force +$\infty$+ ∞, then a terminal NE exists by Lemma 15. If player 2 cannot force +$\infty$+ ∞, then in view of Remark 7, we can consider the SP interdiction game defined by the dual independence system and conclude again by Lemma 15 that it has a terminal NE. The theorem follows.
Acknowledgments
The third and forth authors were working within the framework of the HSE University Basic Research Program. The fourth author was supported in part by the state assignment topic no. FFNG-2024-0003.
Appendix A Proof of Lemma 14
The algorithmic proof is based on the following modification of Dijkstra’s algorithm from [17, 16].
1
Data: A digraph $G=(V,E)$ with arc weights $r:E\to\mathbb{R}_{+}$, a terminal vertex $t\in V$, and a blocking system $\mathcal{I}\in 2^{E}$ defined by an independence oracle.
Result: A potential function $\phi:V\to\mathbb{R}$, and independent sets $\{\widehat{I}(u)\in\mathcal{I}(u)~{}|~{}u\in V\setminus\{t\}\}$, satisfying (7)
2 // Initialization:
3 $S=V\setminus\{t\}$; $T=\{t\}$
4 $\phi(u)=+\infty$ for $u\in S$; $\phi(t)=0$
5 $\widehat{I}(u)=\emptyset$ for $u\in S$
6 Initialize an empty binary heap $\mathcal{H}$
7 for each arc $e=(v,t)\in E$ do
8      Insert $e$ into $\mathcal{H}$ with key $k(e)=r(e)$
9 end for
10
11// Iteration loop:
12 while $\mathcal{H}\neq\emptyset$ do
13       Extract the minimum-key arc $e=(u,v)$ from $\mathcal{H}$
14       if $u\in S$ and $v\in T$ then
15             if $\widehat{I}(u)\cup\{e\}\in\mathcal{I}(u)$ then
16                   $\widehat{I}(u)=\widehat{I}(u)\cup\{e\}$
17            else
18                   $S=S\setminus\{u\}$; $T=T\cup\{u\}$; $\phi(u)=k(e)$
19                   for each arc $e=(v,u)\in E$ such that $v\in S$ do
20                         Insert $e$ with key value $k(e)=r(e)+\phi(u)$ into $\mathcal{H}$
21                   end for
22
23             end if
24
25       end if
26
27 end while
return $\phi$, $\{\widehat{I}(u)~{}|~{}u\in V\setminus\{t\}\}$
Algorithm 1 Modified Dijkstra’s Algorithm $(G=(V,E),t,\mathcal{I},r)$
It easy to see that the running time of the algorithm is $O(|E|(\log|V|+\tau))$. To see that the function $\phi$, and independent sets $\{\widehat{I}(u)\in\mathcal{I}(u)~{}|~{}u\in V\setminus\{t\}\}$ output by the algorithm satisfy (7), let us show that the algorithm maintains the invariant that (7) hold for all $u\in T\setminus\{t\}$. Initially, when $T=\{t\}$, the invariant holds trivially. For the sake of a contradiction, consider the first iteration of the algorithm in which the invariant is violated. Let $e=(u,v)$ be the extracted arc in step 1 in that iteration. Clearly, the invariant remains valid if $u$ is not moved from $S$ to $T$. Assume therefore, that steps 1-1 are executed and let us denote in this case $e=(u,v_{u})$. Then, by the updates used to construct $\widehat{I}(u)$ in step 1 and the condition checked in step 1, we get that
$\widehat{I}(u)\cup\{(u,v_{u})\}\not\in\mathcal{I}(u).$ (11)
By the update in step 1 and the fact that arcs across the $(S,T)$-cut are extracted from $H$ in increasing key-order, we have
$r(u,v_{u})+\phi(v_{u})=\phi(u)\geq r(u,v)+\phi(v),\text{ for all}(u,v)\in\widehat{I}(u).$ (12)
Thus, (7a), (7c) and (7d) hold by (11) and (12). Now consider (7b). For any arc $(u,v)\in E(u)\setminus\widehat{I}(u)$ with $v\in S$, we have $\phi(v)=+\infty$, and therefore $r(u,v)+\phi(v)\geq\phi(u)$ holds trivially. On the other hand, any arc $(u,v)\in E(u)\setminus\widehat{I}(u)$ with $v\in T$ was considered among the arcs crossing the $(S,T)$-cut, prior to moving $u$ to $T$, but was not extracted from $\mathcal{H}$. This implies that $k(u,v)=r(u,v)+\phi(v)\geq k(e)=\phi(u)$. We conclude that (7b) holds for $u$. Let us next consider a vertex $u^{\prime}\in T$ such that $(u^{\prime},u)\in E(u^{\prime})$ and suppose that after updating $\phi(u)$, we get $\phi(u^{\prime})>r(u^{\prime},u)+\phi(u)$. The fact that (7c) holds for all $v\in T\cup\{u\}\setminus\{t\}$ implies that there is a $(u,t)$-path $p=\{(u_{0},u_{1}),\ldots,(u_{k-1},u_{k})\}$ in $G$ such that $u_{0}=u$, $u_{k}=t$ and $\phi(u_{i-1})=r(u_{i-1},u_{i})+\phi(u_{i})$, for $i=1,\ldots,k$. Consider the moment when $u^{\prime}$ was moved from $S$ to $T$; there existed a vertex on $p$, say $u_{j}$, such that the arc $(u_{j},u_{j+1})$ belonged to the $(S,T)$-cut at that moment, but was not extracted from $\mathcal{H}$. This implies that $\phi(u^{\prime})\leq r(u_{j},u_{j+1})+\phi(u_{j+1})$. We get
$r(u^{\prime},u)+r(p[u,u_{j+1}])+\phi(u_{j+1})=r(u^{\prime},u)+\phi(u)<\phi(u^{\prime})\leq r(u_{j},u_{j+1})+\phi(u_{j+1})$
which contradicts the non-negativity of $r$. Finally, note that when $\mathcal{H}$ becomes empty, we have $\phi(u)=+\infty$ for every $u\in S$. Furthermore, all arcs $(u,v)\in E(u)$ with $u\in S$ and $v\in T$ are contained in $\widehat{I}(u)$. This together with the assumption that $E(u)\not\in\mathcal{I}(u)$ imply that any $u\in S$ must have an arc $(u,v_{u})\in S$. Thus, (7) also holds for all $u\in S$ at the end of the algorithm. The lemma follows.